OpenStudy (anonymous):

What is the value of sin 35 ?

OpenStudy (unklerhaukus):

are you looking for the exact value , or an approximation.?

OpenStudy (anonymous):

exactly y

OpenStudy (unklerhaukus):

oh, its just a bit harder

OpenStudy (anonymous):

you want it exactly in surd form?!

OpenStudy (anonymous):

Is it 0.58 ?

OpenStudy (unklerhaukus):

That is an in accurate approximation , to find the exact value use this \[\sin(u \pm v) = \sin (u) \cos (v) ± \cos (u) \sin (v)\]

OpenStudy (unklerhaukus):

\[35=u+v\]

OpenStudy (anonymous):

@ UnkleRhaukus what should I let for u and v for 35?

OpenStudy (unklerhaukus):

well choose so that you can use \(\sin(30°) =1/2 \)

OpenStudy (anonymous):

sin35=cos55=0.573 <----- value when plugged in the calculator. If you want to exact value I strongly suggest you check the unit circle.